### v4

@@ Abstract
In this study, we introduce CT-LLM, a 2B large language model (LLM) that illustrates a pivotal shift towards prioritizing the Chinese language in developing LLMs. Uniquely initiated from scratch, CT-LLM diverges from the conventional methodology by primarily incorporating Chinese textual data, utilizing an extensive corpus of 1,200 billion tokens, including 800 billion Chinese tokens, 300 billion English tokens, and 100 billion code tokens. This strategic composition facilitates the model’s exceptional proficiency in understanding and processing Chinese, a capability further enhanced through alignment techniques. Demonstrating remarkable performance on the CHC-Bench, CT-LLM excels in Chinese language tasks, and showcases its adeptness in English through SFT. This research challenges the prevailing paradigm of training LLMs predominantly on English corpora and then adapting them to other languages, broadening the horizons for LLM training methodologies. By open-sourcing the full process of training a Chinese LLM, including a detailed data processing procedure with the obtained Massive Appropriate Pretraining Chinese Corpus (MAP-CC), a well-chosen multidisciplinary Chinese Hard Case Benchmark (CHC-Bench), and the 2B-size Chinese Tiny LLM (CT-LLM), we aim to foster further exploration and innovation in both academia and industry, paving the way for more inclusive and versatile language models.
@@ Appendix F Training Curves of DPO
The following Figures 6 6 6 6 are the training curves of CT-LLM-SFT-DPO. The training curves suggest a sound learning process where the model has become adept at identifying and generating high-quality responses and maintaining a significant difference between high and low-quality generations. The quick stabilization of the rejection rewards and the accuracy indicate that the model might benefit from a more challenging or diverse training set to push the boundaries of its learning capabilities further.

### v5

@@ Abstract
In this study, we introduce Chinese Tiny LLM (CT-LLM), a 2B large language model (LLM) that illustrates a pivotal shift towards prioritizing the Chinese language in developing LLMs. Uniquely initiated from scratch, CT-LLM diverges from the conventional methodology by primarily incorporating Chinese textual data, utilizing an extensive corpus of 1,200 billion tokens, including 800 billion Chinese tokens, 300 billion English tokens, and 100 billion code tokens. This strategic composition facilitates the model’s exceptional proficiency in understanding and processing Chinese, a capability further enhanced through alignment techniques. Demonstrating remarkable performance on the CHC-Bench, CT-LLM excels in Chinese language tasks, and showcases its adeptness in English through SFT. This research challenges the prevailing paradigm of training LLMs predominantly on English corpora and then adapting them to other languages, broadening the horizons for LLM training methodologies. By open-sourcing the full process of training a Chinese LLM, including a detailed data processing procedure with the obtained Massive Appropriate Pretraining Chinese Corpus (MAP-CC), a well-chosen multidisciplinary Chinese Hard Case Benchmark (CHC-Bench), and the 2B-size CT-LLM, we aim to foster further exploration and innovation in both academia and industry, paving the way for more inclusive and versatile language models.
@@ 6.1 Results of Metrics
6.2 Cultural Biases Evaluation
Figure 3: Political spectrum positioning of CT-LLM compared to other open-source models. We test the models’ orientation with the benchmark introduced by Feng et al. (2023).
The cultural and ideological leanings of pretrained language models (LMs) can provide significant insights into the inherent biases of their training data, as well as the design objectives of their developers. In Figure 3, we observe the distribution of models across different quadrants, particularly distinguishing those trained predominantly on Chinese data from those trained on more global or Western datasets. Our model, CT-LLM, primarily trained on Chinese corpora, is located in the community-focused and structured quadrant. This indicates that the model emphasizes collective cooperation and development, reflecting common themes in Chinese literature and online platforms. In contrast, Western models like TinyLlama-1.1B and Bloom-1.7B, which are based on diverse Western corpora, are situated in the individual-focused and structured quadrant. This suggests that these models highlight personal autonomy and individual expression, focusing on diverse and inclusive approaches.
Interestingly, when examining the openness axis, we find that multilingual models such as MindLLM-1B3-chat-zh and Deepseek-coder-1.3B-instruct exhibit similar cultural characteristics despite being trained across different language corpora. This could be due to the globalized nature of the digital information these models are exposed to. These findings are consistent with observations that LMs exhibit a variety of perspectives and tendencies across different dimensions. Recent web-based texts might reflect a more open stance prevalent in current digital discourse, while those grounded in older corpora might embody more traditional views. To further understand these tendencies, it is necessary to conduct qualitative analysis of the LMs’ responses to various viewpoints. Such analysis would reveal whether these tendencies consistently appear across different discussions, providing a deeper understanding of the presence and nature of any biases in these models.
@@ Appendix F Training Curves of DPO
The following Figures 7 7 7 7 are the training curves of CT-LLM-SFT-DPO. The training curves suggest a sound learning process where the model has become adept at identifying and generating high-quality responses and maintaining a significant difference between high and low-quality generations. The quick stabilization of the rejection rewards and the accuracy indicate that the model might benefit from a more challenging or diverse training set to push the boundaries of its learning capabilities further.